Inverse Laplace Transforms

Find L^-1 {F(s)} when F(s) is given by


1. 1/(s+3)(s+7)


1
Expert's answer
2022-05-25T12:54:21-0400

1.


"\dfrac{1}{(s+3)(s+7)}=\dfrac{A}{s+3}+\dfrac{B}{s+7}"

"=\dfrac{As+7A+Bs+3B}{(s+3)(s+7)}"

"A+B=0"

"7A+3B=1"

"A=\dfrac{1}{4}, B=-\dfrac{1}{4}"


"L^{-1}(\dfrac{1}{(s+3)(s+7)})=\dfrac{1}{4}L^{-1}(\dfrac{1}{s+3})"


"-\dfrac{1}{4}L^{-1}(\dfrac{1}{s+7})=\dfrac{1}{4}e^{-3t}-\dfrac{1}{4}e^{-7t}"